The second-order rate constant has been measured at different temperatures for the decomposition of nitrous oxide (N20) into a nitrogen molecule and an oxygen atom. M T(°C) 1.87 x 10 3600.0 0.0113 .0569 0.244 50.0 00.0 50.0 Determine the activation energy for the reaction. mol
media%2F4d1%2F4d1ac34c-75c4-4a79-927e-47
Answer #1

ANS)

from above data that

now we have to calculate the  ln(k) and 1/T (remembering to convert temperature to kelvin in the process)

bu using above table we can plot the graph between ln ln(k) and 1/T

-6.28 .00115
-4.48 .00108
-2.87 .00103
-1.41 .000978

-2 -3 -4 y = 26.943-29006 x Slope=-E/R -5 -6 -7 0.001 0.00105 0.0011 0.00115 -1 1/T (Kl)

we get a linear graph decreasing ln(k) with increasing 1/T. The slope of this line is equal to -activation energy/R where R is the gas constant.

Using my calculator I found the slope to be -29006

from graph -Ea/R

-29006=-activation energy/R, where R=8.314J/K*mol

Activation energy=2.41*10^5 J/mol

Activation energy=2.41*10^2 kJ/mol

as per HomeworkLib rules we can answered only one question time